Four years ago the age of Paul was twice that of John. John is 18 years old now. How old is Paul now?
Answer by ikleyn(52781)


Four years ago John was  18-4 = 14 years old.

Hence,  Paul was  2*14 = 28 years old four years ago.

Therefore,  Paul is  28+4 = 32 years now.